# Direct multi-model dark-matter search with gravitational-wave interferometers using data from the first part of the fourth LIGO-Virgo-KAGRA observing run

### v1

@@ Abstract
Gravitational-wave detectors can probe the existence of dark matter with exquisite sensitivity. Here, we perform a search for three kinds of dark matter – dilatons (spin-0), dark photons (spin-1) and tensor bosons (spin-2) – using three independent methods on the first part of the most recent data from the fourth observing run of LIGO–Virgo–KAGRA. Each form of dark matter could have interacted with different standard-model particles in the instruments, causing unique differential strains on the interferometers. While we do not find any evidence for a signal, we place the most stringent upper limits to-date on each of these models. For scalars with masses between $4\times 10^{-14},1.5\times 10^{-13}]$ eV that couple to photons or electrons, our constraints improve upon those from the third observing run by one order of magnitude, with the tightest limit of $\sim 10^{-20}\,\text{GeV}^{-1}$ at a mass of $\sim 2\times 10^{-13}\text{ eV}$. For vectors with masses between $7\times 10^{-13},8.47\times 10^{-12}]$ eV that couple to baryons, our constraints supersede those from MICROSCOPE and Eöt-Wash by one to two orders of magnitude, reaching a minimum of $\sim 5\times 10^{-24}$ at a mass of $\sim 10^{-12}$ eV. For tensors with masses of $4\times 10^{-14},8.47\times 10^{-12}]$ eV (the full mass range analyzed) that couple via a Yukawa interaction, our constraints surpass those from fifth-force experiments by four to five orders of magnitude, achieving a limit as low as $\sim 8\times 10^{-9}$ at $\sim 2\times 10^{-13}$ eV. Our results show that gravitational-wave interferometers have become frontiers for new physics and laboratories for direct multi-model dark-matter detection.
@@ I Introduction
The existence of dark matter (DM) is well established, yet its fundamental nature remains one of the most pressing mysteries in modern physics Bertone and Hooper (2018). Across a vast range of possible masses and interaction strengths, dark matter could take many forms – from particles with masses of $\mathcal{O}(10^{-22})$ eV, whose wave-like behavior spans galactic scales Ferreira (2021), to primordial black holes with masses comparable to stars Carr et al. (2016, 2021); Green and Kavanagh (2021). Among these possibilities, ultralight dark matter with masses $m_{\rm DM}\ll 1\,\mathrm{eV}$ presents a particularly compelling scenario. Not only do high-energy theories, including string theory, naturally predict such particles, but their extraordinarily low mass also implies a quantum-mechanical coherence over macroscopic distances, effectively turning them into oscillating background fields. If these fields interact weakly with standard model particles, they could produce distinctive, nearly monochromatic signals detectable by precision experiments Pierce et al. (2018); Grote and Stadnik (2019).
Gravitational-wave (GW) detectors, such as LIGO, Virgo and KAGRA Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021), have already reshaped astrophysics by measuring the signals from merging black holes and neutron stars Abbott et al. (2016, 2017, 2021). But their extraordinary sensitivity also makes them powerful tools for probing fundamental physics beyond GWs Arvanitaki et al. (2015a); Bertone et al. (2020); Nguyen and Miller (2024); Miller (2025). The same precision that captures the faint signals from GWs could reveal subtle interactions between dark matter and ordinary matter, also placing these detectors at the frontier of searches for new physics. Along these lines, multiple searches for dilatons and dark photons have been performed on previous GW data Guo et al. (2019); Abbott et al. (2022); Vermeulen et al. (2021); Fukusumi et al. (2023); Göttel et al. (2024), resulting in stringent constraints on dilatons and dark photons.
In contrast to previous works, we explore here how GW detectors can be used to search for ultralight dark matter in various forms simultaneously: scalar fields (such as dilatons, which may induce oscillations in fundamental constants) Stadnik and Flambaum (2015a, b, 2016); Grote and Stadnik (2019), vector fields (like dark photons, exerting weak oscillating forces on matter) Pierce et al. (2018), and massive tensor fields (closely resembling GWs but with a nonzero mass) Aoki and Mukohyama (2016); Marzola et al. (2018); Armaleo et al. (2021); Manita et al. (2024). Each of these candidates leaves a unique imprint on GW detectors, offering a new way to constrain – or perhaps discover – some of the most elusive dark matter candidates predicted by theory.
@@ II.1 Scalar, dilaton dark matter
Models of ultralight dark matter predict that it could have originated through a vacuum misalignment mechanism in the early universe, and would manifest today as a coherently oscillating scalar field Stadnik and Flambaum (2015b). The coupling of such DM would effectively induce an oscillation of the values of fundamental constants, namely the electron rest mass and the fine structure constant, thus leading to the expansion and contraction of solids. When a field gradient exists, objects can also be accelerated. Many ideas have been proposed to detect such effects Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Kim (2009), with the current strongest direct constraints coming from experiments looking for the oscillation of fundamental constants using atomic clocks or spectroscopy Arvanitaki et al. (2015b); Aharony et al. (2021); Savalle et al. (2021); Kennedy et al. (2020); Antypas et al. (2019, 2021); Tretiak et al. (2022); Oswald et al. (2022); Campbell et al. (2021); Beloy et al. (2021); Zhang et al. (2023); Sherrill et al. (2023). More recently, several coupling paths were discovered to detect those same oscillations using gravitational-wave interferometers Stadnik and Flambaum (2015a, b, 2016); Grote and Stadnik (2019); Morisaki and Suyama (2019); Morisaki et al. (2021). The potential for DM-induced acceleration of the arm mirrors was used to derive upper limits in LIGO data Fukusumi et al. (2023), while oscillations in the size of the beamsplitter was used with the GEO600 detector Vermeulen et al. (2021) and the Fermilab holometer Aiello et al. (2022). An additional coupling path through arm mirror size oscillations was discovered and used in Göttel et al. (2024); Göttel and Raymond (2025), and is used in this study.
Considering a simple linear expansion to the Standard Model Lagrangian $\mathcal{L}$, we can write the scalar field $\phi$ as Vermeulen et al. (2021):
@@
In LIGO, by considering only oscillations in the size of the beamsplitter and test arm mirrors, we can relate the measured strain $h(\omega)$ to the aforementioned DM coupling constants as Göttel et al. (2024):
@@
where $\rho_{\text{DM}}=$0.4\text{\,}\mathrm{GeV}\text{\,}{\mathrm{cm}}^{-3}$\$ is the local DM energy density according to the standard galactic DM halo model Read (2014) and $f_{0}=\omega/(2\pi)$ is the DM frequency. $A_{\rm cal}$, see Appendix A, is obtained through detailed optical simulations that account for finite light travel time effects Morisaki et al. (2021), and encodes the interferometer response to scalar DM Göttel et al. (2024).
@@ II.2 Vector dark photon dark matter
Spin-1 dark photons [^1] These dark photons are different than those that kinetically mix with the ordinary photon Holdom (1986). could completely explain the relic abundance of DM Agrawal et al. (2020a). These particles could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al. (2020b); Co et al. (2019); Bastero-Gil et al. (2019); Dror et al. (2019), or from cosmic string network decays Long and Wang (2019). The observable effect would result from a coupling of dark photons to standard-model particles – either baryons or baryon minus leptons. In particular, this interaction would cause “dark” force on the mirrors, causing them to oscillate at a frequency fixed by the mass of the dark photon Pierce et al. (2018); Miller et al. (2021).
@@
The strain on the interferometers caused by a dark photon DM signal is Pierce et al. (2018); Morisaki et al. (2021):
@@
where $\epsilon_{0}$ is the permittivity of free space, Q/M{Q}/{M} is the charge-to-mass ratio of the mirrors in LIGO and $\omega=2\pi f_{0}$. We take ${Q}/{M}=m_{p}^{-1}$, where $m_{p}=0.93827\,\mathrm{GeV}/c^{2}$ is the proton rest mass Navas et al. (2024).
A second strain also appears because light takes a finite amount of time to travel between the input and end mirrors in the interferometer, during which the mirrors have moved in response to the ultralight DM field Morisaki et al. (2021):
@@ II.3 Tensor dark matter
A novel massive spin-2 particle, derived from a massive gravity theory, could modify gravity and emerge as a promising DM candidate. The foundational framework for linear massive gravity was established by Fierz and Pauli in 1939 Fierz and Pauli (1939). This theory was extended to the nonlinear level, enabling the incorporation of massive gravitons. However, this generalization introduced a non-stable degree of freedom known as the Boulware-Deser ghost Boulware and Deser (1972). To address this issue, a ghost-free massive gravity theory was constructed de Rham and Gabadadze (2010); de Rham et al. (2011), which later evolved into the bimetric gravity Hassan and Rosen (2012) and multigravity theories Hinterbichler and Rosen (2012). In particular, in bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), two spin-2 particles produced by the misalignment mechanism Preskill et al. (1983) – one massless and one massive – could interact and explain DM Marzola et al. (2018).
The Fierz-Pauli Lagrangian density Armaleo et al. (2020) describes this tensor field $\chi_{\mu\nu}$:
@@
where $\chi\equiv\chi^{\mu}_{\mu}$ and $\mathcal{E}_{\mu\nu,\alpha\beta}$ is a second-order derivative operator defined in Eq. 2.13 of Aoki and Mukohyama (2016).
Using the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, one can derive the equations of motion for this ultralight field at late times in the universe Marzola et al. (2018); Aoki and Maeda (2018); Kun et al. (2018).
The strain on interferometers arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha$; thus, the strain can be derived Aoki and Mukohyama (2016); Manita et al. (2024) [^2] Note that Refs. Aoki and Mukohyama (2016); Manita et al. (2024) and Refs. Marzola et al. (2018); Armaleo et al. (2021) differ by a factor of 2 in the strain induced by tensor DM, which results from a factor of 8 difference in the Lagrangians. We use the prescriptions in Aoki and Mukohyama (2016); Manita et al. (2024) throughout this paper, which leads to a factor of two stronger upper limits than would be obtained using the formalism of Marzola et al. (2018); Armaleo et al. (2021):
@@
where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$, $n,m$ are unit vectors pointing down each interferometer arm, $\varepsilon_{ij}$ describes each of the five polarizations of the ultralight DM field Armaleo et al. (2021), $\alpha$ is the Yukawa coupling, and $m_{\text{Pl}}$ is the reduced Planck mass.
@@ III Data
We analyze data from the first part of the fourth observing run (O4a) of the LIGO Livingston (L1) and Hanford (H1) detectors. The sensitivity of the detectors in this observing run improved significantly across all frequencies, particularly above 400 Hz Ganapathy et al. (2023); Jia et al. (2024); Capote et al. (2025). The strain data were collected between May 24, 2023, at 15:00 UTC, and January 16, 2024, at 16:00 UTC, with L1 and H1 operating 69.0% and 67.5% of the time, respectively GWO (2025). Virgo was not operational during O4a, and KAGRA only observed for one month. The data are calibrated by directing a laser onto the test masses and comparing the measured response to the expected one Karki et al. (2016); Viets et al. (2018). Additionally, the data have been cleaned to remove short-duration disturbances (known as “glitches”) Soni et al. (2025), and are only analyzed if they are in “science mode” Goetz and Riles (2025). At worst, amplitude and phase uncertainties at $1\sigma$ are 10% and 10 degrees, respectively, between the frequency range analyzed by all methods: [10, 2000] Hz.
In this analysis, we use three independent pipelines, described in detail in Section IV, to search for direct DM interactions with GW interferometers. Each one employs data in different formats. The first, called the BSD-excess-power method, uses Band Sampled Data (BSD) structures Piccinni et al. (2019) as the input to this search, which are derived from short fast Fourier transform databases Astone et al. (2005). BSDs contain complex-valued time-series representations of the data every $10\,\mathrm{Hz}\,$ every month, and are derived from the strain channel GDS-CALIB_STRAIN_CLEAN_AR. The other two methods, cross-correlation and Logarithmic Power Spectral Density (LPSD), use gated [^3] Gating refers to removing the cumulative effect of many loud, transient glitches that can collectively raise the noise floor and inhibit the recovery of CW signals. data from the channel GDS-CALIB_STRAIN_CLEAN_GATED_G02, in which spurious glitches have been removed. The cross-correlation method employs 1800-s long short Fourier transforms (SFTs) created from G02 gated data during the O4a period for H1 and L1. In total, 5696 pairs of coincident SFTs for H1 and L1 were used, corresponding to a total of 2848 hours. In contrast, the data used by LPSD were divided in continuous segments of at least $1\text{\times}{10}^{5}\text{\,}\mathrm{s}$ each, resulting in 30 segments totaling about $1157\text{\,}\mathrm{h}$, similarly to Göttel et al. (2024).

### v3

@@ Abstract
Gravitational-wave detectors can probe the existence of dark matter with exquisite sensitivity. Here, we perform a search for three kinds of dark matter – dilatons (spin-0), dark photons (spin-1) and tensor bosons (spin-2) – using three independent methods using the first part of the most recent data from the fourth observing run of LIGO–Virgo–KAGRA. Each form of dark matter could have interacted with different standard-model particles in the instruments, causing unique differential strains on the interferometers. While we do not find any evidence for a signal, we place the most stringent upper limits to-date on each of these models. For scalars with masses between $4\times 10^{-14},1.5\times 10^{-13}]$ eV that couple to photons or electrons, our constraints improve upon those from the third observing run by one order of magnitude, with the tightest limit of $\sim 10^{-20}\,\text{GeV}^{-1}$ at a mass of $\sim 2\times 10^{-13}\text{ eV}$. For vectors with masses between $7\times 10^{-13},8.47\times 10^{-12}]$ eV that couple to baryons, our constraints supersede those from MICROSCOPE and Eöt-Wash by one to two orders of magnitude, reaching a minimum of $\sim 5\times 10^{-24}$ at a mass of $\sim 10^{-12}$ eV. For tensors with masses of $4\times 10^{-14},8.47\times 10^{-12}]$ eV (the full mass range analyzed) that couple via a Yukawa interaction, our constraints surpass those from fifth-force experiments by four to five orders of magnitude, achieving a limit as low as $\sim 8\times 10^{-9}$ at $\sim 2\times 10^{-13}$ eV. Our results show that gravitational-wave interferometers have become frontiers for new physics and laboratories for direct multi-model dark-matter detection.
@@ I Introduction
The existence of dark matter (DM) is well established, yet its fundamental nature remains one of the most pressing mysteries in modern physics Bertone:2016nfn. Across a vast range of possible masses and interaction strengths, dark matter could take many forms – from particles with masses of $\mathcal{O}(10^{-22})$ eV, whose wave-like behavior spans galactic scales Ferreira:2020fam, to primordial black holes with masses comparable to stars Carr:2016drx; Carr:2019kxo; Green:2020jor. Among these possibilities, ultralight dark matter with masses $m_{\rm DM}\ll 1\,\mathrm{eV}$ presents a particularly compelling scenario. Not only do high-energy theories, including string theory, naturally predict such particles, but their extraordinarily low mass also implies a quantum-mechanical coherence over macroscopic distances, effectively turning them into oscillating background fields. If these fields interact weakly with standard model particles, they could produce distinctive, nearly monochromatic signals detectable by precision experiments Pierce:2018xmy; Grote:2019uvn.
Gravitational-wave (GW) detectors, such as LIGO, Virgo and KAGRA 2015CQGra..32g4001L; 2015CQGra..32b4001A; KAGRA:2020tym, have already reshaped astrophysics by measuring the signals from merging black holes and neutron stars Abbott:2016blz; gw170817FIRST; LIGOScientific:2021djp. But their extraordinary sensitivity also makes them powerful tools for probing fundamental physics beyond GWs Arvanitaki:2014wva; Bertone:2019irm; Nguyen:2024fpq; Miller:2025yyx. The same precision that captures the faint signals from GWs could reveal subtle interactions between dark matter and ordinary matter, also placing these detectors at the frontier of searches for new physics. Along these lines, multiple searches for dilatons and dark photons have been performed on previous GW data Guo:2019ker; LIGOScientific:2021odm; Vermeulen:2021epa; Fukusumi:2023kqd; Gottel:2024cfj, resulting in stringent constraints on dilatons and dark photons.
In contrast to previous works, we explore here how GW detectors can be used to search for ultralight dark matter in various forms simultaneously: scalar fields (such as dilatons, which may induce oscillations in fundamental constants) Stadnik2015a; Stadnik2015b; Stadnik2016; Grote:2019uvn, vector fields (like dark photons, exerting weak oscillating forces on matter) Pierce:2018xmy, and massive tensor fields (closely resembling GWs but with a nonzero mass) Aoki:2016zgp; Marzola:2017lbt; Armaleo:2020efr; Manita:2023mnc. Each of these candidates leaves a unique imprint on GW detectors, offering a new way to constrain – or perhaps discover – some of the most elusive dark matter candidates predicted by theory.
@@ II.1 Scalar, dilaton dark matter
Models of ultralight dark matter predict that it could have originated through a vacuum misalignment mechanism in the early universe, and would manifest today as a coherently oscillating scalar field Stadnik2015b. The coupling of such DM would effectively induce an oscillation of the values of fundamental constants, namely the electron rest mass and the fine structure constant, thus leading to the expansion and contraction of solids. When a field gradient exists, objects can also be accelerated. Many ideas have been proposed to detect such effects Preskill:1982cy; Abbott:1982af; Dine:1982ah; Cho:2007cy, with the current strongest direct constraints coming from experiments looking for the oscillation of fundamental constants using atomic clocks or spectroscopy Arvanitaki:2014faa; Aharony:2019iad; Savalle:2020vgz; Kennedy:2020bac; Antypas:2019qji; Antypas:2020rtg; Tretiak:2022ndx; Oswald:2021vtc; Campbell:2020fvq; BACON:2020ubh; Zhang:2022ewz; Sherrill:2023zah. More recently, several coupling paths were discovered to detect those same oscillations using gravitational-wave interferometers Stadnik2015a; Stadnik2015b; Stadnik2016; Grote:2019uvn; PhysRevD.100.123512; Morisaki:2020gui. The potential for DM-induced acceleration of the arm mirrors was used to derive upper limits in LIGO data Fukusumi:2023kqd, while oscillations in the size of the beamsplitter was used with the GEO600 detector Vermeulen:2021epa and the Fermilab holometer Aiello:2021wlp. An additional coupling path through arm mirror size oscillations was discovered and used in Gottel:2024cfj; Gottel:2025aqm, and is used in this study.
Considering a simple linear expansion to the Standard Model Lagrangian $\mathcal{L}$, we can write the scalar field $\phi$ as Vermeulen:2021epa:
@@
In LIGO, by considering only oscillations in the size of the beamsplitter and test arm mirrors, we can relate the measured strain $h(\omega)$ to the aforementioned DM coupling constants as Gottel:2024cfj:
@@
where $\rho_{\text{DM}}=$0.4\text{\,}\mathrm{GeV}\text{\,}{\mathrm{cm}}^{-3}$\$ is the local DM energy density according to the standard galactic DM halo model Read:2014qva and $f_{0}=\omega/(2\pi)$ is the DM frequency. $A_{\rm cal}$, see Appendix A, is obtained through detailed optical simulations that account for finite light travel time effects Morisaki:2020gui, and encodes the interferometer response to scalar DM Gottel:2024cfj.
@@ II.2 Vector dark photon dark matter
Spin-1 dark photons [^1] These dark photons are different than those that kinetically mix with the ordinary photon Holdom:1985ag. could completely explain the relic abundance of DM Agrawal:2018vin. These particles could arise from the misalignment mechanism nelson2011dark; arias2012wispy; graham2016vector, parametric resonance or the tachyonic instability of a scalar field agrawal2020relic; Co:2018lka; Bastero-Gil:2018uel; Dror:2018pdh, or from cosmic string network decays long2019dark. The observable effect would result from a coupling of dark photons to standard-model particles – either baryons or baryon minus leptons. In particular, this interaction would cause “dark” force on the mirrors, causing them to oscillate at a frequency fixed by the mass of the dark photon Pierce:2018xmy; Miller:2020vsl.
@@
The strain on the interferometers caused by a dark photon DM signal is Pierce:2018xmy; Morisaki:2020gui:
@@
where $\epsilon_{0}$ is the permittivity of free space, Q/M{Q}/{M} is the charge-to-mass ratio of the mirrors in LIGO and $\omega=2\pi f_{0}$. We take ${Q}/{M}=m_{p}^{-1}$, where $m_{p}=0.93827\,\mathrm{GeV}/c^{2}$ is the proton rest mass ParticleDataGroup:2024cfk.
A second strain also appears because light takes a finite amount of time to travel between the input and end mirrors in the interferometer, during which the mirrors have moved in response to the ultralight DM field Morisaki:2020gui:
@@ II.3 Tensor dark matter
A novel massive spin-2 particle, derived from a massive gravity theory, could modify gravity and emerge as a promising DM candidate. The foundational framework for linear massive gravity was established by Fierz and Pauli in 1939 Fierz:1939ix. This theory was extended to the nonlinear level, enabling the incorporation of massive gravitons. However, this generalization introduced a non-stable degree of freedom known as the Boulware-Deser ghost Boulware:1972yco. To address this issue, a ghost-free massive gravity theory was constructed deRham:2010ik; deRham:2010kj, which later evolved into the bimetric gravity Hassan:2011zd and multigravity theories Hinterbichler:2012cn. In particular, in bimetric gravity Hassan:2011zd; Schmidt-May:2015vnx, two spin-2 particles produced by the misalignment mechanism Preskill:1982cy – one massless and one massive – could interact and explain DM Marzola:2017lbt.
The Fierz-Pauli Lagrangian density Armaleo:2020yml describes this tensor field $\chi_{\mu\nu}$:
@@
where $\chi\equiv\chi^{\mu}_{\mu}$ and $\mathcal{E}_{\mu\nu,\alpha\beta}$ is a second-order derivative operator defined in Eq. 2.13 of Aoki:2016zgp.
Using the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, one can derive the equations of motion for this ultralight field at late times in the universe Marzola:2017lbt; Aoki:2017cnz; Kun:2018ino.
The strain on interferometers arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha$; thus, the strain can be derived Aoki:2016zgp; Manita:2023mnc [^2] Note that Refs. Aoki:2016zgp; Manita:2023mnc and Refs. Marzola:2017lbt; Armaleo:2020efr differ by a factor of 2 in the strain induced by tensor DM, which results from a factor of 8 difference in the Lagrangians. We use the prescriptions in Aoki:2016zgp; Manita:2023mnc throughout this paper, which leads to a factor of two stronger upper limits than would be obtained using the formalism of Marzola:2017lbt; Armaleo:2020efr:
@@
where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$, $n,m$ are unit vectors pointing down each interferometer arm, $\varepsilon_{ij}$ describes each of the five polarizations of the ultralight DM field Armaleo:2020efr, $\alpha$ is the Yukawa coupling, and $m_{\text{Pl}}$ is the reduced Planck mass.
@@ III Data
We analyze data from the first part of the fourth observing run (O4a) of the LIGO Livingston (L1) and Hanford (H1) detectors. The sensitivity of the detectors in this observing run improved significantly across all frequencies, particularly above 400 Hz LIGOO4Detector:2023wmz; membersoftheLIGOScientific:2024elc; Capote:2024rmo. The strain data were collected between May 24, 2023, at 15:$0$ UTC, and January 16, 2024, at 16:$0$ UTC, with L1 and H1 operating 69.0% and 67.5% of the time, respectively GWOSC:O4a. Virgo was not operational during O4a, and KAGRA only observed for one month. The data are calibrated by directing a laser onto the test masses and comparing the measured response to the expected one Karki:2016pht; Viets:2017yvy. Additionally, the data have been cleaned to remove short-duration disturbances (known as “glitches”) LIGO:2024kkz, and are only analyzed if they are in “science mode” GoetzRiles:T2400058:v1. At worst, amplitude and phase uncertainties at $1\sigma$ are 10% and 10 degrees, respectively, between the frequency range analyzed by all methods: [10, 2000] Hz.
In this analysis, we use three independent pipelines, described in detail in Section IV, to search for direct DM interactions with GW interferometers. Each one employs data in different formats. The first, called the BSD-excess-power method, uses Band Sampled Data (BSD) structures Piccinni:2018akm as the input to this search, which are derived from short fast Fourier transform databases Astone:2005fj. BSDs contain complex-valued time-series representations of the data every $10\,\mathrm{Hz}\,$ every month, and are derived from the strain channel GDS-CALIB_STRAIN_CLEAN_AR. The other two methods, cross-correlation and Logarithmic Power Spectral Density (LPSD), use gated [^3] Gating refers to removing the cumulative effect of many loud, transient glitches that can collectively raise the noise floor and inhibit the recovery of CW signals. data from the channel GDS-CALIB_STRAIN_CLEAN_GATED_G02, in which spurious glitches have been removed DavisEtAl2024_selfgating_O4a; DCHgate2025. The cross-correlation method employs 1800-s long short Fourier transforms (SFTs) created from G02 gated data during the O4a period for H1 and L1. In total, 5696 pairs of coincident SFTs for H1 and L1 were used, corresponding to a total of 2848 hours. In contrast, the data used by LPSD were divided in continuous segments of at least $1\text{\times}{10}^{5}\text{\,}\mathrm{s}$ each, resulting in 30 segments totaling about $1157\text{\,}\mathrm{h}$, similarly to Gottel:2024cfj.